The game consists of drawing a card from a deck. You win Php 100.00 if you draw the ace of spades or lose Php 5.00 if you draw any other card.


1
Expert's answer
2021-09-13T00:29:14-0400

"E(X)=\frac{1}{52}*100.00+\frac{51}{52}*(-5)=-\frac{155}{52}\approx-Php2.98."